### v1

# Towards Mobile Sensing with Event Cameras on High-mobility Resource-constrained Devices: A Survey

Abstract
With the increasing complexity of mobile device applications, these devices are evolving toward high mobility. This shift imposes new demands on mobile sensing, particularly in terms of achieving high accuracy and low latency. Event-based vision has emerged as a disruptive paradigm, offering high temporal resolution, low latency, and energy efficiency, making it well-suited for high-accuracy and low-latency sensing tasks on high-mobility platforms. However, the presence of substantial noisy events, the lack of inherent semantic information, and the large data volume pose significant challenges for event-based data processing on resource-constrained mobile devices. This paper surveys the literature over the period 2014-2024, provides a comprehensive overview of event-based mobile sensing systems, covering fundamental principles, event abstraction methods, algorithmic advancements, hardware and software acceleration strategies. We also discuss key applications of event cameras in mobile sensing, including visual odometry, object tracking, optical flow estimation, and 3D reconstruction, while highlighting the challenges associated with event data processing, sensor fusion, and real-time deployment. Furthermore, we outline future research directions, such as improving event camera hardware with advanced optics, leveraging neuromorphic computing for efficient processing, and integrating bio-inspired algorithms to enhance perception. To support ongoing research, we provide an open-source Online Sheet with curated resources and recent developments. We hope this survey serves as a valuable reference, facilitating the adoption of event-based vision across diverse applications.
Index Terms:
Mobile Sensing, Event Camera, Event-based Vision
I Introduction
Figure 1: Mobile devices are employed in a wide range of applications. Among these, the most critical steps involve estimating the devices’ states, perceiving surrounding environmental structures, and understanding the relationships between devices and their environment. As mobile devices advance toward high-mobility design, these mobile platforms impose new demands on mobile sensing, particularly in terms of high accuracy and low latency. This necessitates close coordination between sensors and sensing algorithms: $(i)$ Sensors must acquire high-accuracy data with low latency; $(ii)$ Sensing algorithms must process the data efficiently and accurately under constrained resources. Existing sensors fail to fully meet these requirements; Event cameras, capable of asynchronously capturing pixel-level intensity changes with $\mu s$-level latency, hold the potential to revolutionize mobile sensing. This survey provides a comprehensive review of event cameras and the development of efficient and accurate algorithms tailored to them.
Mobile sensing. With ongoing advancements in sensor technology and the proliferation of sophisticated computing capabilities within embedded systems, mobile devices (e.g., drones and autonomous vehicles) have emerged as the most groundbreaking innovations in recent years [1, 2, 3]. As illustrated in Fig.1, these devices are increasingly deployed in a variety of novel applications, including last-mile delivery [4, 3], industrial inspection [5, 6, 7], rapid relief-and-rescue [8, 9], aerial imaging [10, 11] and sky networking [12, 13], particularly within smart city scenarios. To perform these tasks, which require extensive interaction with the external environment, mobile devices must possess the ability to: $(i)$ awareness their own state, including location and orientation [14, 15, 16], $(ii)$ comprehend their surroundings, (e.g., environmental structure and map) [17, 18], and $(iii)$ understand their relationship with the environment, such as the spatio-temporal relationships between mobile devices and objects within their surroundings [19]. Achieving these capabilities has become a focal point of interest within the mobile computing community.
High mobility trend of mobile devices. In context of smart city scenarios, the demand for mobile devices is progressively increasing, with their mission profiles evolving toward execution of 4D tasks, which are characterized as deep, dull, dangerous, and dirty [20, 21, 22]. The expanding scale of cities also necessitates that mobile devices complete various tasks within shorter time frames, driving their evolution toward high-speed operation [1, 23, 24]. Consequently, the development of mobile devices is exhibiting new trends toward high mobility. For instance, DJI’s industrial inspection drones can cruise at speeds of 21 $m/s$ [25], while Wing’s delivery drones fly at 30 $m/s$ to deliver packages [26].
New challenges for mobile sensing. As mobile devices evolve toward high mobility design, mobile sensing is required to advance toward high accuracy and low latency, enabling these mobile devices to perceive their state and surroundings in $millimeter$-level accuracy with $millisecond$-level latency, thereby facilitating faster responses and more precise adjustments. This evolution establishes new objectives for the sensors and data processing algorithms involved in mobile sensing tasks: $(i)$ On the sensor input front, it is essential to acquire higher accuracy raw data with lower latency. $(ii)$ On data processing algorithms front, efficient processing of sensor measurements is essential to enhance accuracy in mobile sensing tasks while optimizing performance on resource-constrained platforms.
Existing sensors for mobile sensing. However, existing sensors are increasingly inadequate in meeting the high accuracy and low latency requirements for mobile sensing, particularly in environmental sensing and interaction for high-mobility devices. $(i)$ Radar-based solutions leverage sensors such as LiDAR [1, 27], mmWave radar [28, 29], or ultrasound radar [30], which emit signals within specific spectrums and calculate distances from reflected signals. These methods track changes in distances to update device locations and infer spatial relationships with surrounding objects. However, the high sensing delay, low sensing accuracy of these sensors render them poorly suited for high-mobility mobile devices. LiDAR can achieve millimeter-level accuracy, but it requires accumulating points into frames with a low output frequency (e.g., 10 $Hz$), causing a delay equal to the frame interval (typically 100 $ms$). Meanwhile, mmWave radar provides millisecond-level latency but suffers from limited spatial resolution, preventing millimeter-level sensing accuracy. $(ii)$ Camera-based methods utilize monocular [31, 32, 33] and stereo camera [34, 35] for self-state estimation and environmental understanding via SLAM techniques [36]. However, they are computationally intensive and constrained by limited temporal resolution (e.g., <30 $Hz$), high sensing latency (e.g., >30 $ms$), and standard dynamic range (e.g., 60 $dB$), making them unsuitable for high-mobility mobile platforms.
TABLE I: Summary of topics covered in various studies
Topic Design of event camera Advantage of event camera Generation model of event Hardware design Products Datasets Represen-tations Denoising Filtering and feature extraction Matching Mapping Accelera-tion Applica-tion Advantage in mobile compu-ting Challenge in mobile compu-ting
[37]
[38]
[39]
[40]
[41]
This
Survey
New sensor: Event camera. The event cameras are novel bio-inspired sensors that outputs pixel-wise intensity changes in an asynchronous manner [42, 37]. Unlike frame cameras, it generates output based on scene dynamics rather than a global clock that is independent of the scene [43, 44]. The event cameras offer four key advantages that align well with the requirements of mobile sensing tasks for high-mobility mobile devices: $(i)$ The $\mu s$-level temporal resolution enables the capture of high-speed motions without motion blur, supporting accurate sensing during high-speed operations [45, 46]. $(ii)$ The $\mu s$-level sensing latency allows the report of environmental changes to mobile devices almost instantaneously [47]. $(iii)$ The high dynamic range (HDR), which is 140 $dB$ compared to 60 $dB$ for standard cameras, making it effective in diverse lighting conditions [48, 43]. $(iv)$ The low power consumption (e.g., 0.5 $W$) makes it particularly suitable for efficient designed mobile devices [49]. These advantages position event cameras as a promising technology to empower mobile devices designed for high-speed operation and efficiency.
Event-based data processing algorithms. Although event cameras enable high-accuracy data acquisition with low-latency, their unique characteristics introduce three challenges to data processing: $(i)$ Sensitivity to illumination results in significant noise in the sensor output. $(ii)$ The lack of inherent semantic information leads to feature extraction failure. $(iii)$ The data volume is substantial (e.g., thousands of events may be generated within a short time), leading to excessive computation overhead for mobile devices. These challenges make efficient and accurate event-based data processing essential for resource-constrained, high-mobility mobile devices, allowing them to accurately perceive their state and understand their surroundings with minimal latency. This survey categorizes event-based data processing algorithms into six key phases and provides a comprehensive review of each: event representation, event-based denoising, event-based filtering and feature extraction, event-based matching, event-based mapping, and event-based hardware/software acceleration. For specific applications, several end-to-end learning algorithms have been proposed to fully exploit data-driven approaches for task optimization.
Difference between existing surveys. As shown in Fig.2, unlike existing surveys that summarize the applications of event cameras across different domains [37, 38, 39, 40, 41], this survey focuses on how event cameras aid high-mobility resource-constrained mobile devices in high-accuracy and low-latency self-state estimation and environmental understanding, surveying the literature over the period 2014-2024. It outlines processing workflow of event data and reviews the latest advancements at each stage of this workflow. Using the key metrics of accuracy and efficiency in mobile computing, we summarize various methods in each stage to provide a deeper analysis of cutting-edge research. It also covers works on event-based hardware and software acceleration, offering critical insights for deploying event cameras on resource-constrained mobile devices. Finally, the survey discusses application scenarios of event cameras on mobile platforms.
Figure 2: Structure of this survey.
Contribution. The main contribution of the survey paper is summarized as follows.
$\bullet$ We present a comprehensive review of how event cameras enhance high-mobility resource-constrained mobile devices by enabling high-accuracy, low-latency self-state estimation and environmental perception.
$\bullet$ We provide a comprehensive introduction to event generation models, event camera hardware design, commercial products, as well as event-based datasets.
$\bullet$ We highlight unique advantages of event cameras in mobile sensing, as well as the specific challenges they encounter.
$\bullet$ We categorize event stream processing methods into several stages and provide a comprehensive review of each, including event stream representation, data processing algorithms, acceleration techniques, and applications on event-based mobile platforms.
$\bullet$ We present our insights and solutions for future trends, with a particular focus on bio-inspired event camera hardware design, algorithm development, and hardware-software co-optimization techniques.
Online resource. This survey presents a comprehensive review of event-based sensing systems, focusing on key technological advancements and practical applications. To further support the research community, we have established a open-source Online Sheet  [^1] Event-based mobile sensing resource that will be regularly updated, ensuring access to the latest developments and fostering continued innovation in event-based sensing system.
Organization. Fig.2 illustrates the structure of this survey. In Sec. II, we introduce the primer of event cameras, including the principles behind event generation, the hardware design of event cameras, existing products, and datasets. We also emphasize the advantages and challenges of applying event cameras to mobile devices. Sec. III discusses methods for representing event data, while Sec. IV reviews existing work in event data processing, including event-based denoising, filtering, matching, and mapping. This survey highlights the performance of current methods in terms of accuracy and efficiency. In Sec. V, we present hardware and software acceleration techniques that enable event cameras to operate effectively on resource-constrained mobile devices. Sec. VI covers the application scenarios of event cameras in mobile devices. In Sec. VII, we propose potential future research directions, followed by a conclusion in Sec. VIII.
II Primer: Event camera development
Figure 3: Principle of the Event Cameras: Events are generated based on changes in logarithmic light intensity over time.
II-A Event generation model
Unlike conventional cameras that capture images at fixed intervals, event cameras operate asynchronously by detecting changes in log intensity at individual pixels, generating events only when significant changes occur. This characteristic enables event cameras to achieve exceptionally high temporal resolution and effectively mitigate motion blur, particularly in scenarios involving fast-moving objects.
Each event is defined by the pixel location where the change occurs, the timestamp, and the polarity. Formally, an event can be represented as:
$e_{k}=(x_{k},t_{k},p_{k}),$
where $e_{k}$ denotes the event, $x_{k}$ specifies the pixel location, $t_{k}$ represents the timestamp, and $p_{k}$ indicates the polarity of the change. The change in log intensity is given by:
$\Delta L(x_{k},t_{k})=L(x_{k},t_{k})-L(x_{k},t_{k}-\Delta t_{k}),$
where $L(x_{k},t_{k})$ represents the log intensity at pixel $x_{k}$ and time $t_{k}$. As shown in Fig.3, an event is triggered only when $|\Delta L(x_{k},t_{k})|$ exceeds a predefined threshold $C$. The polarity $p_{k}$ is assigned as $+1$ if $\Delta L(x_{k},t_{k})>0$, and $-1$ otherwise. This event-driven paradigm substantially reduces redundant data, enhancing processing efficiency, conserving computational resources, and enabling deployment in resource-constrained systems such as embedded devices [50, 51].
Figure 4: Working Mechanism of Event Cameras: Inspired by the rod cells in the human eye, event camera operates at the pixel level, independently transforming light into voltage signals to capture intensity variations.
In practical applications, the negative threshold $C^{-}$ and positive threshold $C^{+}$ need not be identical and can be tuned to meet specific requirements. These thresholds play a critical role in determining the performance of event cameras. If the threshold $C$ is set too high, the camera may overlook subtle changes, thereby reducing its sensitivity. Conversely, if $C$ is set too low, noise may trigger an excessive number of events, leading to redundant event generation. Hence, careful selection of threshold values is essential for achieving optimal performance in event-based systems.
When the time interval $\Delta t_{k}$ is sufficiently small, the change in log intensity $\Delta L(x_{k},t_{k})$ can be approximated using a Taylor series expansion, yielding:
$\Delta L(x_{k},t_{k})=\frac{\partial L(x_{k},t_{k})}{\partial x}\Delta t_{k}.$
By substituting $\Delta L(x_{k},t_{k})$ with $p_{k}C$, the expression becomes:
$\frac{\partial L(x_{k},t_{k})}{\partial x}=\frac{p_{k}C}{\Delta t_{k}},$
which provides an indirect method for estimating the gradient of intensity. Assuming uniform illumination over the interval $\Delta t_{k}$, the following relationship can be derived:
$\Delta L=-\nabla L\cdot v\Delta t_{k},$
where $v$ represents the velocity of moving edges. This formula indicates that events are generated as a result of edge motion.
TABLE II: Classification of Event Camera Types
Indicator Original Event Camera [52] DVS [53] ATIS [54] DAVIS [55]
Development Origin Mahowald, Mead at Caltech Inspired by silicon retina Evolved from DVS Advanced DVS/ATIS
Output Data Type Log brightness Change in brightness Change + absolute brightness Change + absolute brightness
Sensor Type Large pixels, CMOS CMOS CMOS CMOS
Pixel Structure Single-pixel design Smaller, simpler pixels Dual-subpixel Shared pixel/subpixel
Brightness Measurement Continuous-time Change only Absolute + change Absolute + change
Dynamic Range Limited Narrow High Moderate
Event Synchronization Basic Fast reset Potential mismatch Slow sampling
Noise Filtering Minimal Simple Complex Advanced
II-B Hardware design
The first event camera was developed by Mahowald and Mead at Caltech, producing data in the form of logarithmic brightness changes [52]. As the precursor to modern event cameras, it exhibited several limitations, such as the requirement for precise bias potentiometer adjustments, large pixel dimensions, and restricted practical applicability. In this part, we will introduce hardware design of modern event cameras, as illustrated in Tab. II.
General hardware architecture. As illustrated in Fig.4, event cameras typically employ CMOS sensors due to their ability to achieve low latency. The working principle of CMOS sensors can be summarized as follows: $(i)$ When incident light interacts with the CMOS sensor, photons are absorbed by the semiconductor material, generating electron-hole pairs. The quantity of these pairs is directly proportional to the intensity of the incident light. $(ii)$ The electrons are subsequently drawn to the storage region under the influence of an applied electric field. $(iii)$ Readout circuits then convert the accumulated electrons into a voltage signal corresponding to the light intensity. This voltage signal is used to calculate changes in logarithmic brightness.
After the CMOS sensor detects a change in logarithmic brightness, the data is processed by the event generation circuit. This circuit applies noise filtering before comparing the filtered data to a predefined threshold using a comparator. If the brightness change exceeds the threshold, the circuit generates an event, which is promptly transmitted to the processor via a serial communication module.
To ensure temporal accuracy, event cameras incorporate a time synchronization circuit that employs high-precision synchronization protocols to maintain consistent timestamps for events. In scenarios involving large data volumes that cannot be processed in real-time, event cameras utilize temporary memory for data buffering. This mechanism ensures data integrity, even under high-throughput conditions.
DVS event camera [53]. The design of the Dynamic Vision Sensor (DVS) draws inspiration from the frame-based silicon retina architecture. It employs a circuit to detect changes in brightness from a continuous-time photoreceptor, leveraging capacitance coupling for this purpose. After each measurement, the circuit resets its state to prepare for subsequent readings. The DVS exclusively outputs changes in brightness rather than absolute brightness values. One notable advantage of the DVS is its relatively simple circuit architecture, which enables smaller pixel sizes compared to other event cameras. However, its output is limited to event data, posing challenges in extracting sufficient information for more advanced processing tasks (e.g., static scenarios).
ATIS event camera [54]. The pixels in an Active Time-Image Sensor (ATIS) are composed of subpixels, with one subpixel dedicated to measuring absolute brightness. This design results in each ATIS pixel having an area approximately twice that of a DVS pixel. By incorporating this feature, ATIS achieves both a wide dynamic range and a broad static range. The dynamic range refers to the ratio between the strongest and weakest signals detectable by the sensor, with a high dynamic range ensuring image clarity under extreme lighting conditions, whether very bright or very dark. Nevertheless, a key limitation of ATIS is the potential mismatch between the absolute brightness measurements and event data, particularly in high-speed motion scenarios. This discrepancy arises because absolute brightness is computed as an average across all pixels, whereas events are triggered at individual pixels. Consequently, when significant changes occur in localized regions of the image, the absolute brightness may fail to capture these variations accurately, as such changes exert minimal influence on the global brightness average.
DAVIS event camera [55]. The Dynamic and Active Vision Sensor (DAVIS) is capable of outputting both absolute brightness and event-based data. In DAVIS, pixels and subpixels share the same sensor, enabling a more compact design. As a result, the pixel area is smaller than that of the ATIS, with only a modest 5% increase in size compared to the DVS. However, the sampling speed of the DAVIS circuit is slower than that of the DVS circuit. This reduced sampling speed can lead to temporal mismatches between the absolute brightness measurements and event data, potentially causing motion blur, especially in high-speed scenarios.
TABLE III: Specifications of Different Event Camera Models
Supplier iniVation Prophesee
Camera model DVXplorer DVXplorer Micro DAVIS346 DAVIS346 AER EVK4 HD EVK5 HD
Indicator Resolution (pixels) 640 $\times$ 480 640 $\times$ 480 346 $\times$ 260 346 $\times$ 260 1280 $\times$ 720 1280 $\times$ 720
Latency (µs) 1000 1000 <1000 <1 <100 <800
Dynamic range (dB) 90-120 110 120 120 >86 >110
Pixel Size (µm) 9 9 18.5 18.5 4.86 x 4.86 4.86 x 4.86
Power Consumption Max 12W <140 mA @ 5 VDC Typical 180 mA @ 5 VDC Typical 180 mA @ 5 VDC Typical 0.5W Typical 0.5W
Sync Support Multi-camera sync Multi-camera sync No multi-camera sync No multi-camera sync External trigger (Type B) External trigger (4 Pin)
Temporal Resolution (µs) 200 200 1 1 100 100
Physical diagram [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
II-C Advantages of event cameras in mobile sensing
Compared to frame cameras, event cameras offer numerous potential advantages for mobile devices. They have the potential to significantly enhance the sensing capabilities of mobile devices, improve operational efficiency, and expand their range of applications. Specifically, the advantages that event cameras provide to mobile devices include:
High temporal resolution. The high mobility of mobile devices leads to rapid environmental changes, causing motion blur in standard frame cameras and making timely sensing about changes in environment and self-state difficult for radar-based or camera-based solution due to their low spatial-temporal resolution. Event cameras, with microsecond resolution and the ability to detect changes without motion blur, offer crucial support for quickly sensing changes in the environment and device state.
Low latency. The high mobility of mobile devices requires rapid awareness of environmental and self-state changes to respond effectively. Standard frame cameras depend on a global exposure time (about 20 $ms$) and additional processing time (10-20 $ms$), which delays reaction times and increases the risk of collisions. In contrast, event cameras operate with independent pixels that generate events immediately upon brightness changes, without requiring a global exposure time. This results in sub-millisecond latency, allowing mobile devices to detect changes and respond quickly.
High dynamic range (HDR). Mobile devices are increasingly used in challenging environments, such as low-light nighttime and bright daytime settings, requiring reliable sensing across varying lighting conditions. Standard frame cameras have a dynamic range of about 60 $dB$, making them less effective in extreme lighting. In contrast, event cameras operate on a logarithmic scale with independently functioning pixels, offering a very high dynamic range with >120 $dB$. This enables them to adapt to both extremely dark and bright conditions, making mobile devices suitable for a wider range of scenarios.
Low power consumption. Mobile devices, designed for operation in complex urban environments, often prioritize efficiency, which can limit their computational capabilities and power reserves. Standard frame cameras typically demand substantial computational resources to process image data, leading to increased processing times and additional power consumption for mobile devices. In contrast, event cameras transmit only changes in brightness, thereby eliminating redundant data. This elimination in data redundancy alleviates the computational and power burdens on mobile devices, enhancing their operational efficiency.
In summary, event cameras offer numerous advantages for mobile devices, expanding their application scenarios and enhancing efficiency. At the same time, mobile devices can leverage benefits of event cameras to unlock full potential.
II-D Challenges of event cameras in mobile sensing
Since event cameras operate fundamentally differently from frame-based cameras by capturing per-pixel brightness changes asynchronously as events, their integration into mobile devices poses several challenges:
(i) How to mitigate event bursts and accurately extract features from event data, given high-speed operation of mobile devices and its lack of inherent semantic information? Event cameras are highly sensitive to illumination changes, with even minor variations triggering numerous events. On high-mobility mobile devices, rapid scene changes captured by onboard event cameras can trigger event bursts, generating thousands of events in a short time. Event bursts and the lack of inherent semantic information hinder the extraction of critical environmental information, as they are often overwhelmed by irrelevant events caused by device motion.
(ii) How to efficiently process large volume of event data given on-board constrained resources? Mobile devices typically employ low-power embedded systems for efficiency, which inherently limits computational resources. Meanwhile, rapid scene changes captured by onboard event cameras produce large volumes of events in a short time, requiring efficient processing to support high-speed operation. This demand places significant strain on the limited computational capacity, compromising efficiency and hindering the full utilization of event cameras’ potential.
II-E Commercial product & comparation
II-E1 Commercial companies and products
Event cameras are increasingly entering the commercial market, with several leading companies spearheading the development and production of these advanced sensors. Notable manufacturers include Prophesee, Inivation, and Celepixel, each providing specialized products tailored to diverse application scenarios. These companies are at the forefront of innovation in mobile sensing areas such as robotics, autonomous vehicles, industrial automation, and machine vision, capitalizing on the benefits of event-based sensing technologies to enable real-time, high-speed performance in demanding applications.
Inivation [56]. Inivation, a prominent leader in the event camera industry, specializes in developing high-resolution, energy-efficient event cameras. Their flagship sensors, such as the DAVIS240 and DAVIS346, uniquely combine event-based and traditional frame-based data, providing a valuable platform for researchers to analyze and compare these two modalities in real-time. Inivation’s cameras offer resolutions of up to 1 megapixel (1 MP), positioning them among the highest-resolution event cameras available. Designed for exceptional low-light performance and durability in outdoor environments, these sensors are well-suited for diverse applications, including robotics and autonomous systems.
Prophesee [57]. Prophesee is a leading innovator in event-based vision technology, renowned for its cameras’ ultra-high dynamic range exceeding 120 dB. This makes them particularly well-suited for environments with variable lighting conditions, such as autonomous driving and industrial automation. One of Prophesee’s flagship products, the Metavision sensor and EVK4 HD and EVK5 HD event cameras based on this sensor, features an exceptional temporal resolution in the microsecond range, enabling it to capture fast-moving objects with minimal latency. This high temporal precision facilitates accurate real-time monitoring and processing, providing significant advantages in applications that demand rapid response times and high accuracy.
II-E2 Comparation of products
Tab. III provides a comparative analysis of various event cameras based on key specifications, including resolution, latency, dynamic range, pixel size, power consumption, interface type, weight, casing material, synchronization support, and additional features. Resolutions range from $346\times 260$ to $1280\times 720$, with latencies spanning from 1 $\mu s$ to 800 $\mu s$, reflecting significant differences in processing speed. Dynamic ranges vary between 86 $dB$ and 120 $dB$, accommodating diverse lighting conditions, while pixel sizes range from 4.86 $\mu m$ to 18.5 $\mu m$, affecting sensitivity and image detail. Power consumption differs notably, with EVK models typically consuming around 0.5 $W$, whereas the DVXplorer can draw up to 12 $W$. Most cameras utilize USB 3.0 interfaces, although some support Type-C or micro-B connectors. Weights vary from 16 $g$ to 120 $g$, and most cameras feature aluminum casings, with the DVXplorer Micro employing POM plastic. Synchronization support is present in most models, except for the DAVIS346 and DAVIS346 AER. Additional features include integrated inertial measurement units (IMUs), neuromorphic compatibility, and adjustable sensor configurations, catering to a wide range of applications.
TABLE IV: Event-based Datasets
Dataset name Year Data Volume Perspective Participants Lighting Conditions Annotation Count Application Scenario
MVSEC [58] 2018 - Dynamic Pedestrians, vehicles Daytime - Driving, handheld scenes
DVS-Pedestrian [59] 2019 0.1 hours, 4.6K annotations Dynamic Pedestrians Daytime 4.6K Walking street
Gen1 [60] 2020 39 hours, 255K annotations Dynamic Pedestrians, vehicles Daytime, night 255K Driving
DDD20 [61] 2020 51 hours Dynamic Pedestrians, vehicles Daytime, night - Driving
1 Megapixel [62] 2020 15 hours, 25M annotations Dynamic Pedestrians, vehicles Daytime, night 25M Driving
DSEC [63] 2021 1 hour, 390K annotations Dynamic Pedestrians, vehicles, scenes Daytime, night 390K Driving
PEDRo[64] 2023 0.5 hours, 43K annotations Dynamic Pedestrians Daytime, night 43K Robotics detection
eTraM [65] 2024 10 hours, 2M annotations Static Vehicles, pedestrians, micro-mobility Daytime, night, twilight 2M Intersections, roadways, streets
LLE-VOS[66] 2024 70 video clips Dynamic Pedestrians, other targets Normal, low-light 5600 Gym, classroom, zoo
LLE-DAVIS[66] 2024 90 video clips Dynamic Various scene objects Standard light, low-light 6118 Synthetic dataset
Synthetic eSfP [67] 2024 90 video clips Static Spinning objects Standard light, low-light - Object rotation
II-F Event-based datasets
The event-based datasets collectively serve as important benchmarks in the field of autonomous driving, robotics, and visual perception (Tab. IV). They provide diverse sensor data, including event cameras, LiDAR, and IMUs, enabling researchers to develop and test algorithms that are capable of handling dynamic, high-speed, and low-light environments. By using these datasets, developers can improve the accuracy, efficiency, and robustness of their algorithms for mobile sensing tasks (e.g., localization, object detection, and SLAM).
MVSEC [58]. The MVSEC (Multi Vehicle Stereo Event Camera) dataset integrates data from event cameras, traditional frame-based cameras, and LiDAR sensors, making it a comprehensive resource for tasks such as stereo vision, SLAM, and autonomous driving. It provides synchronized stereo event streams captured from vehicles navigating diverse driving environments, including urban roads and highways. The dataset’s real-world scenarios and multimodal sensor data make it invaluable for testing and evaluating event camera algorithms in dynamic and complex conditions typical of autonomous driving applications.
DVS-Pedestrian [59]. The DVS-Pedestrian dataset is a benchmark dataset designed specifically for pedestrian detection and tracking using event-based vision. It consists of high-speed, high-resolution event data captured by a DVS as pedestrians move in various urban environments. The dataset includes sequences where pedestrians exhibit diverse behaviors such as walking, running, or standing still, captured in different lighting conditions and backgrounds. Due to the event-based nature of the sensor, it excels in capturing fast-moving objects in challenging scenarios, including low-light and high-speed motion. This makes the DVS-Pedestrian dataset particularly useful for evaluating pedestrian detection algorithms that leverage the unique advantages of event cameras, such as high temporal resolution and low latency.
Gen1 [60]. The Gen1 dataset is a synthetic dataset designed for the development and evaluation of event-based vision systems. It includes high-resolution event camera data, captured in various dynamic scenarios that simulate real-world environments. The dataset provides both monocular and stereo event streams, offering temporal and spatial information with high precision. The synthetic nature of Gen1 allows for controlled conditions and a wide range of variations, such as changes in lighting, motion speed, and scene complexity. Gen1 is particularly useful for evaluating event-based algorithms in tasks like motion estimation, depth estimation, and visual odometry, where real-time performance and accuracy in dynamic environments are critical.
DDD20 [61]. The DDD20 (Dynamic Driving Dataset 2020) is a comprehensive dataset designed for the development and testing of autonomous driving systems. It includes synchronized sensor data from stereo cameras, LiDAR, GPS, and IMU, collected in real-world driving environments. The dataset captures dynamic urban and highway scenes, featuring a wide range of road users such as vehicles, pedestrians, cyclists, and various road obstacles. With its high-resolution video, point cloud data, and detailed ground-truth annotations, DDD20 is an invaluable resource for tasks like object detection, tracking, semantic segmentation, and motion prediction. The diverse and challenging driving scenarios make it an ideal benchmark for evaluating autonomous driving algorithms in realistic, dynamic conditions.
DSEC [63]. The DSEC (Dynamic and Static Environment for Cars) dataset is a large-scale dataset for autonomous driving research, designed to test algorithms in both dynamic and static environments. It contains data from multiple sensors, including event cameras, LiDAR, and RGB cameras, captured from vehicles moving through urban streets and highways. The DSEC dataset is particularly useful for tasks like visual odometry, 3D reconstruction, and object tracking in dynamic environments. Its multimodal nature allows for the development of algorithms that can handle challenges of autonomous driving, such as dealing with fast-moving objects and varying light conditions. This dataset is an important resource for advancing autonomous driving technologies, providing real-world data for complex driving scenarios.
PEDRo [64]. The PEDRo (Pedestrian Detection and Recognition) dataset is a large-scale dataset designed for the evaluation of pedestrian detection and recognition algorithms. It contains a wide variety of real-world urban driving scenarios, with annotated images and event-based data captured from multiple sensors, including monocular and stereo cameras, along with LiDAR and GPS information. The dataset includes diverse pedestrian behaviors, such as walking, running, and interactions with other objects in complex traffic environments. With its rich annotations, including bounding boxes and activity labels, PEDRO is ideal for testing algorithms for pedestrian detection, tracking, and behavior recognition, particularly in the context of autonomous driving systems that need to understand and predict pedestrian movements in dynamic environments.
eTraM [65]. The eTraM (Event-based Traffic Monitoring) dataset is a comprehensive collection of event-based data designed for the development and evaluation of traffic monitoring and analysis systems using event cameras. It includes synchronized event camera data alongside traditional frame-based camera images, LiDAR, and GPS/IMU information, captured in urban road settings. The dataset features dynamic scenes with various types of vehicles, pedestrians, and complex traffic scenarios, providing high temporal resolution and spatial precision. eTraM is particularly useful for tasks such as real-time traffic flow analysis, vehicle tracking, and anomaly detection, offering a rich resource for researchers working on event-based perception and event-driven processing in autonomous driving applications.
LLE-VOS [66]. The LLE-VOS (Low-Light Event-based Video Object Segmentation) dataset is a specialized dataset designed for video object segmentation tasks in low-light environments using event cameras. It provides synchronized event-based data and traditional frame-based video, captured under challenging lighting conditions, including night-time and indoor scenes. The dataset features various dynamic objects, such as moving vehicles, pedestrians, and other environmental elements, with ground-truth object masks for segmentation tasks. The LLE-VOS dataset is particularly useful for developing and evaluating video object segmentation algorithms that can work robustly in low-light and high-motion scenarios. This makes it an important resource for applications in autonomous driving, surveillance, and robotics, where visibility in poor lighting conditions is crucial.
VECtor [68]. The VECtor dataset provides a high-quality, large-scale collection of sensor data from autonomous driving vehicles, specifically focused on high-precision localization and environmental understanding. The dataset includes data from event cameras, LiDAR, and RGB sensors, capturing a wide range of driving environments from urban to highway scenarios. Its primary application is in the development of autonomous driving technologies, particularly for tasks like path planning, object detection, and localization. The high-precision sensor data makes it a valuable resource for testing algorithms that require fine-grained understanding of the environment, such as real-time mapping and localization in highly dynamic settings.
TUM-VIE [69]. The TUM-VIE (TUM Visual-Inertial Evaluation) dataset is designed for evaluating VIO and SLAM algorithms. It contains synchronized data from monocular cameras and IMUs, recorded during various motion scenarios, including both indoor and outdoor environments. This dataset is particularly useful for testing algorithms that combine visual and inertial data to estimate the camera’s position and orientation in real time. Its focus on dynamic environments and the integration of visual and inertial data makes it a valuable resource for developing algorithms that can operate effectively in complex, real-world situations. The TUM-VIE dataset is widely used in research related to robotics and autonomous navigation.
Event-based datasets serve diverse applications, including traffic monitoring, low-light video object segmentation (VOS), autonomous driving, and robotics. Datasets such as eTraM and Gen1 offer data for both static and dynamic traffic monitoring, covering daytime and nighttime conditions. For VOS tasks in low-light environments, datasets like LLE-VOS and LLE-DAVIS combine event and traditional frame data to improve object visibility. Autonomous driving datasets, including MVSEC and DSEC, provide multimodal data by integrating event cameras with LiDAR and RGB sensors, supporting complex tasks like visual odometry and depth estimation. These datasets are essential for advancing event-based vision in real-world and challenging conditions. These datasets collectively demonstrate the strengths of event cameras, particularly in high-speed, low-light, and high dynamic range environments. They play a crucial role in advancing research across various fields, including intelligent transportation systems, robotics, where traditional imaging technologies often face limitations.
Figure 5: Event representations method (a) Raw event (b) 2D histogram (c) Time surface (d) Voxel grid (e) RGB picture.
III Abstraction: Event Representation
Event data is often processed and transformed into various representations to extract meaningful information ("features") for solving specific tasks. Here, we review popular representations of event data, which range from simple, hand-crafted transformations to more elaborate methods Fig.5.
III-A Raw events
Individual Events. Raw events $e_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})$ are utilized by event-by-event processing methods such as probabilistic filters and Spiking Neural Networks (SNNs). These methods build additional information from past events or external knowledge and fuse it with incoming events asynchronously to produce an output. Examples include [70] [71] [72] [73] and [74].
Spiking Neural Networks (SNNs) are the neuromorphic counterpart to conventional neural networks, or Artificial Neural Networks (ANNs). While ANNs, particularly Deep Neural Networks (DNNs), are known for their accuracy, they often neglect computational complexity and energy consumption. SNNs excel in processing event camera data, which captures scene changes asynchronously, producing sparse, spatio-temporal data. By mimicking biological neural networks, SNNs integrate input spikes and fire when a threshold is reached. This spike-based processing is fault-tolerant and energy-efficient, only activating when necessary, unlike the continuous computation in ANNs.
Event Packet. The event set $E=\{e_{k}\}^{N_{e}}$ retains precise timestamp and polarity information for each event. Selecting the appropriate packet size $N_{e}$ is crucial to meet the assumptions of the algorithm (e.g., constant motion speed throughout packet’s duration), which varies depending on the task. Examples can be found in [75] [76] [77] [78] and [79]
Raw events offer high fidelity, retaining complete temporal and spatial information, making them ideal for event-driven processing, especially in applications using SNNs. These networks are energy-efficient and well-suited for real-time tasks due to their spike-based processing, which mimics biological neural networks and only activates when necessary. However, the large volume of data generated by raw events can make direct processing complex, requiring advanced time management and synchronization mechanisms to handle the asynchronous nature of the data effectively. While raw events provide detailed and precise information, they come with the challenge of managing substantial data loads and ensuring the proper alignment of events across time.
III-B Event frame (2D Grid)
Events in a spatio-temporal neighborhood are converted into an image (2D grid) through simple methods such as counting events or accumulating polarity pixel-wise. This image can then be fed into image-based computer vision algorithms.
Histogram. This representation converts events into a histogram, providing a natural activity-driven sample rate. Although this practice is not ideal in the event-based paradigm, it has a significant impact in [80] [81].While traditional 2D histograms capture spatiotemporal event data by discretizing the event stream into bins and counting occurrences over time, an AEIM (Activity-Aware Event Integration Module) enhances traditional 2D histograms by integrating complex spatiotemporal operations and attention mechanisms to capture finer scene details and prioritize high-confidence information, which reduces noise and improves performance in dynamic environments, making it especially effective for tasks of semantic segmentation [82].
Time Surface. A Time Surface (TS) is a 2D map where each pixel stores a single time value (e.g., the timestamp of the last event at that pixel). Events are converted into an image whose "intensity" is a function of the motion history at that location, with larger values corresponding to more recent motion. TSs explicitly expose the rich temporal information of the events and can be updated asynchronously. Examples include [83] [84] [85] [86] [87].
The Event Frame, a 2D grid representation, has several advantages, such as its compatibility with existing image processing algorithms, which makes integration into established systems relatively straightforward. Its simplicity also facilitates easy implementation and understanding, supporting use of well-established computer vision techniques. However, the Event Frame comes with notable disadvantages. It can lose important temporal information, potentially leading to motion blur in dynamic scenes. Additionally, it is not ideal for HDR scenes, where drastic contrasts between bright and dark areas can cause issues. Furthermore, it may not fully exploit the sparse nature of event data, which can lead to inefficiencies in data representation and suboptimal performance in certain scenarios.
III-C Spatio-temporal 3D grid representation
A Voxel Grid is a space-time (3D) histogram of events, where each voxel represents a particular pixel and time interval. This representation preserves the temporal information of the events better by avoiding collapsing them onto a 2D grid. If polarity is used, the voxel grid becomes an intuitive discretization of a scalar field (polarity or brightness variation) defined on the image plane, with the absence of events marked by zero polarity. Each event’s polarity may be accumulated on a voxel or spread among its closest voxels using a kernel, providing sub-voxel accuracy. Examples include [88, 89, 90, 91, 92, 93, 94, 95] .
The Voxel Grid representation offers several advantages, such as retaining detailed temporal and spatial information, which helps to preserve the dynamics of the scene more effectively than simpler 2D representations. It is particularly suitable for 3D processing and complex scene analysis, as it allows for precise localization and tracking of events in space and time. This enables more accurate handling of dynamic scenes compared to traditional approaches. However, this method also has its disadvantages. It requires significant memory and computational resources, making it more demanding than 2D representations. Additionally, the implementation and management of voxel grids are more complex, and the increased data volume may present challenges in real-time processing. To fully exploit the rich information contained in a voxel grid, sophisticated algorithms are required, adding to the complexity of its use.
III-D Customized representation
Customized representations are specifically designed for particular tasks or applications, where traditional representations might not fully capture desired information or where the task requires a more tailored approach. These representations can combine multiple techniques (e.g., temporal and spatial features) or apply domain-specific knowledge to improve the efficiency and accuracy of event-based processing.
For example, in the context of motion deblurring, a customized event representation may incorporate motion models or sensor-specific characteristics to better handle motion blur in the reconstructed images. In [96], the authors propose a customized event representation for motion deblurring that integrates adaptive filtering techniques to preserve high-frequency details, resulting in sharper images. Similarly, in [97], the 2D-1T Event Cloud Sequence (2D-1T ECS) provides a customized representation that effectively separates spatial and temporal components of the event data, enabling the capture of both geometric and motion information. By organizing events into spatial clouds (2D) and temporal sequences (1T), this representation preserves the sparsity of the event data, ensuring efficient processing. Additionally, the inclusion of surface-event-based sampling and local event normalization improves the encoding of spatial and temporal features, making it particularly well-suited for tasks like object classification and action recognition. This tailored approach allows the model to leverage the unique structure of event data while minimizing computational costs.
Task-specific representations offer several benefits, including being optimized for particular tasks or domains, which can lead to better performance. By combining multiple event representations, these models can leverage different features to enhance overall accuracy. Additionally, incorporating domain-specific knowledge helps improve task performance and precision. However, such representations also have their drawbacks. They can be complex to design and implement, requiring careful consideration of the task’s specific needs. Moreover, they may not be easily generalizable across different tasks or domains, limiting their flexibility. If not carefully crafted, these representations could also introduce biases, which may affect the accuracy of the results.
IV Algorithm: Event processing
IV-A Event-based denoising
Motivation. With the increasing adoption of event cameras in applications requiring high-speed processing and low-latency performance, effective event denoising has emerged as a critical step in the field of event-based vision. Event cameras, inspired by biological vision systems, operate using a continuous differential sampling mechanism that detects brightness changes with exceptional temporal resolution. However, this high sensitivity also makes them prone to various noise sources, which can significantly degrade the quality of event streams and impede performance in downstream tasks such as reconstruction and object detection [98].
Noise in event cameras arises from both external and internal sources. For example, ambient light fluctuations and junction leakage currents in the imaging circuitry can generate spurious events even without actual intensity changes, a phenomenon referred to as background activity (BA) [53, 99]. BA not only undermines algorithmic accuracy but also consumes communication bandwidth and computational resources. In high-speed scenarios or during prolonged operation, hot pixels—pixels that fail to reset properly [100]—can further contribute to noise by continuously producing erroneous events, known as hot noise.
Unlike conventional cameras, which benefit from the smoothing effects of image integration, event cameras inherently amplify noise through their logarithmic encoding and differential sampling mechanisms [101]. This amplification exacerbates issues such as BA, false negatives (missed events), and stochastic timing disturbances. Moreover, event cameras often generate redundant trailing events following edge arrivals, adding complexity to the denoising process and presenting further challenges for effective noise mitigation.
Challenge. Event cameras, while promising for high-speed and low-latency applications, face significant challenges in noise mitigation due to the diverse and complex nature of noise, limitations in data annotation, computational efficiency constraints, and non-uniform distribution of noise patterns.
One major challenge lies in the diversity of noise types intrinsic to event cameras: (i) BA noise, triggered by junction leakage currents or low-light conditions, generates spurious events unrelated to actual intensity changes. (ii) Hot pixels, common during prolonged usage or high-speed scenarios, produce persistent erroneous signals. (iii) Temporal noise introduces stochastic variations in event timing, while (iv) structural noise arises from edge inconsistencies or redundant trailing events. The varied spatial, temporal, and statistical properties of these noise sources make their simultaneous mitigation particularly challenging.
The lack of high-quality annotated datasets further hinders the development of effective denoising algorithms. Constructing paired datasets for supervised learning is labor-intensive and often impractical. Although synthetic datasets are commonly used, they fail to capture the complexity and diversity of real-world noise, resulting in significant domain gaps. Multi-modal approaches that integrate data from frames or IMUs offer partial solutions but are limited in challenging conditions such as extreme motion blur or poor illumination.
Efficiency constraints present another critical obstacle, especially for real-time applications like robotics and autonomous navigation. Many denoising methods prioritize accuracy at the cost of computational efficiency, making them unsuitable for latency-sensitive scenarios or power-constrained devices. This challenge becomes even more pronounced in low-light environments, where noise levels surge, further increasing computational demands.
Lastly, the nonuniform spatial and temporal distributions of noise reduce the efficacy of traditional denoising techniques, which often assume uniform noise characteristics [98]. For instance, BA noise frequently exhibits localized bursts, while other disturbances vary dynamically over time. Robust denoising methods must account for irregularities to ensure reliable performance across diverse environments.
[…]
Figure 6: Comparison of different event-based denoising algorithms in computational cost (Gflops) vs. performance (PSNR).
Literature review. Event denoising has achieved notable advancements through diverse methodologies, including traditional signal processing, statistical approaches, machine learning, surface fitting techniques, and biologically inspired frameworks. These developments have significantly improved noise mitigation in event-based systems Fig.6.
Statistical methods. Early event-denoising techniques primarily employed statistical methods to identify and remove outliers in event streams. These methods typically assess event density within a local spatio-temporal neighborhood, classifying low-density events as noise [108]. Delbruck et al. [109] pioneered density-based filtering, leveraging local context to suppress noisy events. Subsequent efforts [110, 103, 104] enhanced these techniques with optimized event storage strategies to reduce computational complexity and improve processing efficiency. However, these approaches often require manual parameter tuning to adapt to varying noise conditions, limiting their scalability and generalizability.
Filtering-based methods. To better exploit the temporal and asynchronous nature of event data, several filtering-based algorithms have been developed, including temporal, spatial, and spatio-temporal filters. (i) Temporal filters [111] leverage the temporal correlation of events generated by object edges to eliminate redundant or ambiguous events. (ii) Spatial filters [112] focus on pixel intensity changes to isolate events related to moving objects. (iii) Spatio-temporal filtering methods [102, 108, 113, 114] combine these strategies to effectively suppress BA noise caused by intensity-independent events. For instance, Liu et al. [110] demonstrated the advantages of integrating spatial and temporal filtering to reduce BA noise while preserving critical motion-related events.
Surface fitting techniques. Surface fitting methods offer an alternative approach, particularly effective for smoothing event data in continuous motion scenarios. Techniques such as EV-Gait [105] and the Guided Event Filter (GEF) [115] utilize local plane fitting, optical flow estimation, and image gradients to filter noisy events. The time-surface (TS) method [86, 111] transforms event streams into a monotonically decreasing representation to address sparsity issues. These methods excel in scenarios with single-object motion but face challenges in low-light or highly dynamic environments.
Deep learning-based methods. Recent advances in deep learning have revolutionized event denoising, enabling automated and adaptive solutions. Deep learning models train on noisy and clean event pairs to learn mappings, reducing reliance on manual parameter tuning. Baldwin et al. [106] proposed the Event Denoising Convolutional Neural Network (EDnCNN), which integrates frames and IMU data to classify events as signal or noise. EventZoom [116] utilizes a U-Net architecture for efficient noise-to-noise denoising, achieving superior performance in handling noisy event streams. AEDNet [107] directly processes raw DVS data, preserving inherent spatio-temporal correlations while denoising. Additional methods, such as K-SVD [117] and multilayer perceptron-based denoising filters (MLPF) [104], focus on sparse feature extraction and event probability estimation, further advancing the capabilities of learning-based denoising frameworks.
These diverse methods collectively try to address the challenges of event denoising, paving the way for more robust and efficient event-based vision systems. However, further research is needed to develop universally adaptable solutions capable of handling diverse noise characteristics and operational constraints.
TABLE V: Comprehensive Comparison of event filtering and feature extraction methods
Method Advantages Disadvantages Applications
Frame-based methods [118, 119] High computational efficiency; Low memory consumption; Simple implementation
Poor accuracy in high-speed scenarios;
Limited temporal precision;
Loss of fine event details
Basic feature detection;
Initial event processing
Surface of Active Events-based methods [120, 121] High temporal accuracy; Precise event timing preservation; Good feature localization
High memory overhead;
Reduced processing efficiency;
Complex computational requirements
Temporal-spatial feature detection;
Optimization tasks.
Clustering-based methods [122, 123] Linear time complexity; Efficient memory usage; Fast stream processing
Accuracy depends on parameters;
Poor precision in complex scenes;
Unstable performance
Dense event processing;
Dynamic scene analysis
Geometric Transform-based methods [124, 125, 126] High tracking accuracy; Precise motion estimation; Robust feature detection
Heavy computational load;
Low processing efficiency;
High resource consumption
Fast motion tracking;
Extreme illumination scenarios
Temporal Filtering-based methods [127, 75] Fast processing speed; Efficient memory utilization; Good real-time performance
Accuracy affected by noise;
Precision loss in filtering;
Detail preservation issues
High temporal precision tasks;
Real-time processing;
Resource-constrained systems
Asynchronous methods [128, 129, 130] Low latency processing; High temporal accuracy; Efficient event handling
Complex implementation;
Resource intensive
Low-latency applications;
High-speed corner detection;
Noise-heavy environments
Hybrid methods [131, 132, 133] High detection accuracy; Robust feature extraction; Multi-layer filtering
Computationally heavy;
Complex parameter tuning
High-speed tasks;
Multi-feature scenarios
Neural Networks-based methods [134, 135, 127, 136] High feature accuracy; Tracks complex dynamics; Biologically inspired
Training-intensive;
High computational demands;
Accuracy-speed tradeoffs
Complex dynamic environments;
Biological vision systems
Frame-Event hybrid methods [137, 138, 139] High spatial-temporal accuracy; Precise feature matching; Robust performance
Complex design;
High computational cost;
Complex resource management
Robotics;
Precision tasks;
High-speed tracking
IV-B Event-based filtering and feature extraction
Motivation. Unlike traditional frame-based systems that capture images at fixed intervals, event cameras generate data asynchronously, triggered only by significant changes in brightness. This results in a sparse and highly dynamic output, necessitating specialized filtering techniques to isolate and retain the most meaningful events for downstream analysis, thereby enabling effective feature extraction.
The primary objective of event filtering is to emphasize events that capture critical scene dynamics—such as motion and edges—while discarding irrelevant or redundant data. This selective approach not only enhances the efficiency of algorithms for tasks like object detection, motion tracking, and scene reconstruction but also reduces computational overhead. In fast-moving or highly dynamic environments, filtering ensures that only the most informative events are processed, improving both the accuracy and speed of downstream tasks. Moreover, effective filtering also plays a pivotal role in mitigating noise inherent in event data, which can otherwise compromise feature extraction for applications like object recognition and scene understanding. This is particularly crucial in scenarios involving rapid motion or abrupt lighting changes, where noise can obscure essential information. By prioritizing relevant data, filtering ensures that critical events are preserved, facilitating more robust and accurate processing for real-time applications.
Challenge. The asynchronous nature of event data introduces significant challenges in defining consistent filtering criteria, as determining what qualifies as a "significant" event often lacks a clear, universal standard. Additionally, the high variability in event frequency and timing complicates the preservation of temporal coherence during feature extraction, which is critical for accurately representing motion. Balancing computational efficiency with complexity of filtering and feature extraction algorithms further exacerbates these challenges; while sophisticated methods may improve filtering quality, they can also introduce processing delays, undermining real-time performance.
Parameter tuning presents another major obstacle, as achieving optimal filter and extraction performance often requires finely adjusted settings that vary across different operational scenarios. This adaptability is especially crucial in real-world applications, where event cameras must contend with highly dynamic and unpredictable environments. Therefore, event filtering plays a pivotal role in enhancing the utility of event-based vision systems, addressing these challenges is essential for advancing feature extraction techniques and ensuring robust performance in dynamic conditions.
Literature review. Event-based feature extraction has evolved significantly, transitioning from adaptations of conventional frame-based methods to specialized event-driven approaches (Tab. V). Early work extended traditional techniques, such as adapting Harris corner detection [118] to binary frames derived from event accumulation [119]. A key breakthrough came with the introduction of the Surface of Active Events (SAE) [120], a representation that stores the temporal information of the most recent events at each pixel. This innovation bridged the gap between frame-based and event-driven paradigms, enabling algorithms to preserve the temporal precision inherent in event cameras while improving computational efficiency and feature extraction performance.
The evolution of extraction methodologies has seen notable improvements in both accuracy and efficiency. The development of eFAST [128] marked a transition from computationally intensive gradient-based methods to faster comparison-based operations optimized for event data. Subsequent methods, such as the Arc* algorithm [121], enhanced detection speed and corner repeatability by refining SAE filtering techniques. Hybrid approaches, including FA-Harris [131] and TLF-Harris [132], integrated efficient candidate selection with robust multi-layer filtering, balancing computational complexity and accuracy. FEAST [134] introduced an unsupervised feature extraction approach using spiking neuron-like units with individual selection thresholds, while ROT-Harris [133] leveraged tree-based processing to enhance feature extraction beyond traditional 2D methods. These advancements reflect the field’s shift toward practical and scalable implementations.
Parallel to advancements in feature extraction, event filtering has emerged as a critical preprocessing step for robust feature extraction, leveraging the asynchronous and high-resolution temporal data of event cameras. Temporal filtering-based methods has significantly improved data quality and computational efficiency by prioritizing meaningful scene changes [127, 75]. Clustering-based methods, such as eCDT [122], dynamically group events to represent dense streams compactly while minimizing complexity. However, such methods often face challenges in maintaining robustness under varying conditions due to sensitivity to parameter configurations [123].
Parametric filtering has further advanced event-based vision by employing geometric transformations to filter irrelevant data. For instance, EKLT [124] aligns events to improve feature extraction, while curve-fitting techniques construct smooth spatio-temporal trajectories [125, 126], excelling in scenarios with rapid motion or extreme illumination. Asynchronous methods, such as HASTE [129], process individual events in real time using hypothesis-driven transformations, and proximity-based trackers continuously refine feature locations, effectively suppressing noise and supporting low-latency applications [121, 130].
The integration of filtering and feature extraction methods has significantly enhanced the robustness and efficiency of event-based vision systems. For example, combining filtering mechanisms with classical algorithms, such as Harris or FAST, has enabled robust corner detection, while shape detection methods using iterative closest point (ICP) and Hough transforms have improved performance in high-speed and high-dynamic-range scenarios [137]. Modern tracking techniques leverage the low latency and high temporal resolution of event cameras by employing probability-based associations and spatio-temporal constraints, further enhancing feature stability and tracking accuracy [138, 139].
Neural network-based filtering represents a promising frontier, combining asynchronous event processing with biologically inspired architectures. Pulse-based neural networks [135], for example, exploit high temporal resolution for precise feature tracking, demonstrating strong potential in dynamic and complex environments. However, current neural network architectures still face challenges in achieving detection accuracy comparable to traditional convolutional models [127, 136]. The continued integration of filtering, feature extraction, and tracking methodologies is driving progress, expanding the application scope of event-based vision systems while maintaining computational efficiency and robustness in diverse real-world scenarios.
IV-C Event-based matching
Motivation. Matching refers to the identification of corresponding features between two or more event streams captured at different times or from different viewpoints. These correspondences serve as a foundation for tasks such as visual odometry, video interpolation, and other event-based mobile sensing applications, enhancing accuracy and robustness in high-frequency dynamic environments. Unlike RGB cameras, which capture images at fixed intervals, event cameras respond only to brightness changes, making them particularly well-suited for dynamic scenes. They excel at capturing edges and texture information within the field of view while avoiding the redundant data accumulation typical of frame-based systems. Additionally, the temporal information embedded in event data enables the inference of object motion. The unique characteristics of event cameras effectively mitigate challenges faced by traditional cameras, including motion blur and low frame rates. Consequently, event cameras offer inherent advantages for matching-based tasks, with even single-modal event data demonstrating strong performance in these applications.
Challenge. Event-based matching tasks, however, present several significant challenges. First, the sparsity of event data in the pixel space and its non-uniform temporal distribution complicate the extraction of sufficient and robust features. This issue is particularly pronounced in low-light conditions or static scenes, where fewer events are generated, making matching significantly more difficult. These scenarios demand the development of specialized feature extraction algorithms tailored to the unique characteristics of event cameras for event-based matching. Moreover, the ultra-high temporal resolution of event cameras generates a substantial volume of event data, leading to a significantly higher computational load compared to traditional frame-based cameras. Addressing this challenge requires not only improving the efficiency of matching algorithms but also ensuring high accuracy with minimal latency to meet the real-time demands of mobile and resource-constrained computing applications.
TABLE VI: Event-based Matching Algorithm Comparison
Algorithm Sample Input Advantages Disadvantages
Local Feature matching [140, 141, 142, 143, 144, 145, 146, 147, 148, 149] Optical Flow Batch Low latency, noise resistance Low computational resource Difficult to handle global motion Difficult to handle repeat texture Need further extraction method
Feature Descriptor: SIFT/SURF/ORB Difficult to handle global motion
Optimization-based matching [150, 151, 152] Contrast Maximization Batch/Asynchronous Global consistency Higher accuracy after iterations Detailed design of optimizer Higher computational complexity Higher time latency Hard to converge Sensitive to initial values
Levenberg-Marquardt
Graph optimization
Neural Network-based matching [153, 154, 155, 156, 44, 157] CNN Batch/Asynchronous Adaptive learning Less manual intervention Efficient inference Large amounts of training data Need parameter tuning Poor generalization ability
SNN
GNN
Literature review. In visual data processing, matching algorithms play a pivotal role in tasks such as motion estimation, visual odometry, and feature tracking. These algorithms can be broadly classified into three categories: local feature matching, optimization-based matching, and deep learning-based matching, each offering distinct approaches, strengths, and limitations Tab. VI.
Local feature matching methods. These methods emphasize detecting and associating local specific features within visual data[158, 159, 160, 161]. Techniques such as optical flow [140, 141, 142, 143] and feature descriptors like SIFT [144], SURF [145], and ORB [146, 147] exemplify this category. These methods are characterized by low latency, robustness to noise, and minimal computational requirements, making them particularly well-suited for resource-constrained applications such as mobile devices, autonomous driving, and robotic vision [148, 149]. However, local feature matching struggles with event stream and challenges like large-scale global motion, repeated textures, and a reliance on specific feature extraction methods [162, 163]. Performance degrades in scenarios with sparsely distributed features or drastic scene changes, also highlighting the need for more advanced techniques to enhance robustness and accuracy.
Optimization-based matching methods. These methods aim to achieve global consistency by solving mathematical optimization problems to align event data or trajectories. Popular techniques include contrast maximization [150, 151, 152], the Levenberg-Marquardt algorithm, and graph-based optimization. These methods are highly flexible and capable of achieving high accuracy through batch processing or asynchronous updates, making them valuable in event camera research. However, they are computationally intensive and sensitive to initial parameter settings. Poor initialization can lead to non-convergence or convergence to local optima. Additionally, the high computational demands and memory requirements make these methods less suitable for real-time applications, especially when processing large-scale data. Balancing computational efficiency and matching accuracy remains a significant challenge in this domain.
Deep learning-based matching methods. These methods have emerged as a transformative paradigm, leveraging advanced models such as Convolutional Neural Networks (CNNs) [155, 156, 44], Spiking Neural Networks (SNNs) [153, 154], and Graph Neural Networks (GNNs) [157]. These methods automatically learn feature representations and matching rules from large-scale training data, surpassing traditional hand-crafted feature descriptors in adaptability and precision. Graph-based approaches, particularly GNNs, effectively exploit the asynchronous nature of event cameras, delivering strong performance in high-frequency dynamic scenarios. Despite their ability to process large volumes of event data efficiently, deep learning methods are computationally demanding and require substantial hardware resources. Additionally, their reliance on extensive labeled training data and their limited generalization in unseen or structurally diverse environments pose challenges. Training deep models is time-consuming, and achieving optimal performance often necessitates significant expertise in hyperparameter tuning.
Selecting the most suitable matching algorithm depends on the specific task requirements, including computational constraints, real-time processing needs, data characteristics, and scene complexity. As technology advances, these algorithms are likely to evolve and converge, enabling the development of more versatile and robust visual matching techniques capable of addressing a broader spectrum of applications.
TABLE VII: Comparison of different mapping methods.
Types of Mapping Methods Advantages Disadvantages or Challenges Frame-based mapping [164, 165, 166, 167, 168, 169] ✓ Improves the accuracy of depth estimation ✓ Enhances the quality of map construction ✓ Lower latency and computation cost ✗ Requires measuring and updating depth information between current and initial detection frames, potentially increasing computational load Filter-based mapping [170, 171, 172, 173] ✓ High robustness, suitable for rapidly changing environments ✓ Dynamically adjusts the system’s map representation ✗ Requires a large number of parameters to represent camera poses, potentially increasing computational complexity    Continuous-time mapping [174, 175, 176] ✓ Reduces parameter complexity ✓ Improves mapping accuracy and efficiency ✓ Simultaneously updates camera poses and 3D landmarks ✗ Requires handling continuous curve interpolation and optimization problems ✗ Increasing computational difficulty ✗ High latency due to frequent states update Spatio-Temporal Consistency[177] ✓ Improves the accuracy of map construction ✓ Optimizes motion parameters ✗ Requires iteratively searching for the closest points and applying a pruned ICP algorithm ✗ High computational cost and latency
IV-D Event-based mapping
Motivation Building on event matching technology, we can obtain a sequence of continuous event streams captured by an event camera from different viewpoints and time instances within its field of view. As the camera pose changes, these events enable a progressive sensing and reconstruction of the surrounding spatial environment. By leveraging a series of camera poses, depth information from each event can be incrementally accumulated and fused, ultimately forming a complete, dense 3D map. In event-based sensing and mapping applications, real-time performance, adaptability, and computational efficiency are key requirements. The spatial sparsity and high temporal resolution of event cameras reduce the computational burden, allowing mapping algorithms to prioritize regions of change while conserving resources. Moreover, the high temporal resolution minimizes motion blur, enabling the camera to capture object edges with precision as it moves through the scene. This not only ensures real-time mapping but also reduces the need for extensive data filtering, providing a robust foundation for constructing precise and efficient event-based maps.
Challenge. Using cameras, including event cameras, for mapping and depth estimation is feasible but presents several challenges. Monocular cameras inherently lack depth information when capturing 3D structures, often necessitating additional processing to infer depth. One approach is leveraging perspective changes across multiple frames, while another involves integrating auxiliary sensor data, such as from an IMU. Alternatively, a stereo event camera system can directly estimate depth through disparity calculations. The high temporal resolution of event cameras introduces significant computational demands, as processing their continuous, asynchronous data stream requires substantial computational resources to maintain real-time performance. Furthermore, aligning event data with other modalities poses additional challenges, particularly in feature extraction and synchronization calibration. These issues become more complex in dynamic environments, where ensuring consistency across multiple sensors is crucial for accurate depth estimation.
Literature review. Mapping plays a fundamental role in constructing a 3D representation of the environment based on visual features captured by cameras. With advancements in event-based vision, event cameras have become an increasingly valuable source of information in visual simultaneous localization and mapping, offering high-frequency, asynchronous data well-suited for real-time processing in dynamic scenes. Various event-based mapping approaches have been proposed, each exhibiting distinct characteristics and advantages Tab. VII.
Frame-based mapping methods. Frame-based mapping methods typically rely on event-based 2D image representations for tracking and map construction. These approaches often incorporate depth filters to estimate scene depth, updating depth values and their associated uncertainties through feature triangulation. Specifically, by refining depth measurements between the current frame and the initial frame, these methods significantly enhance depth estimation accuracy. For instance, the studies in [164] and [165] estimate camera poses relative to a known planar structure by minimizing reprojection error. This involves computing the reprojection error between reprojected lines and corresponding event data, as well as between intersection points of these lines and their associated 3D points. Similarly, [166] and [167] employ the SVO algorithm [168] to estimate relative camera poses by minimizing reprojection error over a set of event feature correspondences. Within the mapping framework, these methods utilize depth filters that model depth values as a combination of Gaussian and uniform distributions, refining them iteratively through feature triangulation. This process compares the current feature location with its initial detection, guided by estimated relative camera poses. Furthermore, [169] formulates a linear system to encode the algebraic distance between 2D and 3D vertical lines, solving for camera poses using least squares optimization to improve mapping precision.
Filter-based mapping methods. Filter-based mapping techniques are designed to accommodate the asynchronous nature of event data, dynamically updating the map representation during camera tracking. A common approach in this category is line-based vSLAM, where the map is updated by measuring the distance between events and reprojected lines, constructing a point cloud representation. These methods extract 3D line features using the Hough transform, leveraging implicit spatial correlations between events. Filter-based approaches have demonstrated high robustness, particularly in environments with rapid scene changes. For example, [170] estimates camera poses by measuring the distance between the back-projected ray of event data and 3D planar feature points. Extending this concept, [171] introduces a measurement function that probabilistically models event occurrences on planar surfaces for improved mapping accuracy. Furthermore, line-based mapping methods such as [172] update filter states during camera tracking by assessing event-to-line distances. These methods often employ Event-Based Multi-View Stereo (EMVS) [173] for point-based map construction and utilize the Hough transform for 3D line extraction, reinforcing event-line associations for enhanced spatial understanding.
Continuous-Time mapping methods. To mitigate the high parameter count associated with discrete camera pose representations in filter-based approaches, continuous-time mapping methods have been introduced. These techniques replace discrete pose estimates with continuous trajectory representations, such as B-splines or Gaussian process motion models, enabling smooth trajectory interpolation from local control states. This not only reduces parameter complexity but also facilitates the joint optimization of camera poses and 3D landmarks, thereby improving mapping accuracy and efficiency. For example, methods proposed in [174, 174] model camera trajectories using B-splines, while [175, 176] employ Gaussian process motion models. By interpolating camera poses at arbitrary time instances, these approaches optimize control states for enhanced mapping performance. Notably, [175] integrates incremental Structure from Motion (SfM) into a joint optimization framework, simultaneously updating control states and 3D landmarks for improved mapping consistency. Additionally, [177] introduces a spatio-temporal consistency constraint, assuming that event pairs with identical time intervals undergo the same relative rotational transformation. By leveraging these constraints, motion parameters can be optimized to enhance mapping accuracy, particularly during rotational camera motion. In practice, spatio-temporal consistency methods iteratively search for the closest corresponding points and apply a pruned Iterative Closest Point (ICP) algorithm to enforce spatial consistency, further improving the precision of event-based map reconstruction.
V Acceleration
Event-based sensing significantly reduces data generation compared to conventional frame-based cameras by producing pulse events solely in response to pixel-level dynamic light intensity changes. This data sparsity offers substantial advantages in computational efficiency, low power consumption, and low-latency processing, positioning event-based vision as a promising foundation for high-speed, energy-efficient solutions in time-sensitive applications.
With the growing demand for mobile and edge computing, event-based vision systems are increasingly deployed on resource-constrained platforms (e.g., autonomous vehicles, drones, and wearable devices). These platforms require real-time, energy-efficient processing to ensure reliable operation in dynamic environments. However, achieving low-latency inference while maintaining energy efficiency presents significant challenges, particularly in mobile computing scenarios where computational resources are constrained, and real-time performance is imperative. Striking a balance between accuracy and computational efficiency is thus crucial for the practical deployment of these systems.
To address these challenges, the development of optimized hardware acceleration platforms and specialized software algorithms is essential. These platforms must fully leverage the sparsity of event-driven data to minimize computational overhead while preserving high accuracy, as shown in Fig.7 and Fig.8. By enabling real-time processing at the mobile platforms, such advancements will unlock the full potential of event-based vision for autonomous systems and other next-generation applications.
V-A Hardware acceleration
Motivation. Event-based vision represents a paradigm shift in visual data acquisition and processing, where only changes in scene dynamics generate data, leading to significantly lower data rates than conventional frame-based systems. The inherent sparsity of event data, coupled with its high temporal resolution and low latency, presents substantial advantages for computational efficiency, particularly in real-time, power-constrained environments. However, fully leveraging these benefits requires specialized hardware platforms capable of effectively managing the asynchronous and sparse nature of event streams, a challenge for conventional processors designed for synchronous dense-frame processing.
In the event-based vision processing pipeline, hardware acceleration plays a pivotal role in bridging the gap between raw event streams and high-level vision tasks such as object detection, tracking, and scene reconstruction. The primary objective is to process asynchronous event streams efficiently while preserving the low-latency and low-power characteristics of event-based cameras. This entails converting sparse input data into structured, usable formats suitable for deep neural networks or other vision algorithms.
Hardware acceleration ensures the efficient processing of event data by leveraging specialized architectures that exploit the inherent sparsity and parallelism of event streams. By optimizing data flow, minimizing redundant computations, and enabling parallel processing pathways, hardware accelerators establish the foundation for high-precision, low-latency and energy-efficient event-based vision systems, facilitating their deployment in real-world applications.
Challenge. Traditional Central Processing Unit (CPU)-based computing is inadequate for real-time processing due to its inherently sequential nature and limited capability to handle the sparse, asynchronous data generated by event cameras. While CPUs offer flexibility as general-purpose processors, they fail to meet the stringent requirements of high-throughput, low-latency applications, such as object detection and tracking, particularly on resource-constrained platforms. Although Graphics Processing Units (GPUs) are widely used to accelerate computations, especially in deep learning applications, they present notable challenges in event-based vision. Designed for dense, parallel computations, GPUs excel in conventional frame-based vision tasks where image data is structured and contiguous. However, their architecture is poorly suited for event-driven inputs, which exhibit high spatial and temporal sparsity. This leads to suboptimal resource utilization, increased latency, and excessive power consumption.
By contrast, Field-Programmable Gate Arrays (FPGAs) present a compelling hardware solution for event-based vision processing due to their reconfigurable and massively parallel architectures. Operating on a reconfigurable hardware paradigm, FPGAs enable customized dataflow architectures tailored to the unique characteristics of event camera data. Their ability to exploit event data sparsity through sparse convolutions and asynchronous processing allows for low-latency, and energy-efficient performance. Unlike GPUs, FPGAs circumvent the need for dense matrix operations, optimizing computation for the sparse and event-driven nature of the input. Additionally, their inherent parallelism and fine-grained control over hardware resources make them particularly well-suited for mobile deployments, where balancing power efficiency and computational performance is essential.
Figure 7: Power consumption of different platforms($W$) vs. Latency(Real Computing time)($ms$)
[…]
Literature review. Recent research on hardware acceleration for event-based vision can be broadly categorized into three main approaches: neuromorphic computing, event-driven deep neural network (DNN) acceleration, and hardware optimization techniques aimed at enhancing efficiency and reducing power consumption.
[…]
Event-driven DNN acceleration. While neuromorphic computing presents a bio-inspired approach, another key research direction focuses on hardware acceleration for event-driven deep learning models. Unlike traditional GPUs, which struggle to exploit the sparsity of event-based inputs, specialized architectures have been developed to enhance efficiency. (i) Sparse dataflow architectures: The combinable dynamic sparse dataflow architecture (ESDA) [183] introduces a configurable dynamic sparse dataflow model on FPGAs. By employing a unified sparse token-feature interface, ESDA interconnects parameterizable network modules, reducing latency and power consumption in event-based DNN inference, making it highly suitable for edge computing. (ii) Optimized fusion for event-based vision. EventBoost [186] accelerates event-image fusion tasks using a dedicated hardware accelerator on Zynq’s commercial SoC platform. By optimizing real-time fusion workloads, EventBoost mitigates the inefficiencies of CPU- and GPU-based processing, significantly improving performance for tasks requiring event-driven visual representations.
Hardware optimization for efficient event processing. Beyond the direct implementation of neuromorphic computing and event-driven DNNs, hardware acceleration is crucial for unlocking the full potential of event-based vision, particularly in applications constrained by power budgets and demanding high efficiency. A critical aspect of optimization involves minimizing power consumption right at the sensor interface. For example, [195] pioneered this approach by introducing a dedicated on-chip DVS camera interface. This innovative interface aggregates asynchronous event streams into ternary event frames, achieving a significant reduction in power expenditure for subsequent processing stages. In addition to front-end optimizations, a range of general-purpose hardware solutions has emerged to further enhance processing efficiency and reduce power consumption across the entire event-based vision pipeline.
[…]
Application-specific hardware acceleration design. The development of specialized hardware architectures adapted to accelerate specific tasks has become a key focus in event-based vision research. By optimizing hardware resources to meet the precise computational demands of targeted applications, these solutions achieve significant performance within their intended domains. A notable example is [179], which introduced an energy-efficient processor-on-chip specifically designed to accelerate hand gesture recognition tasks.
A prominent application domain that is increasingly driving the need for specialized event-based vision hardware is unmanned aerial vehicles (UAVs), or drones, where stringent constraints on size, weight, and power consumption necessitate highly optimized hardware solutions. The systems proposed in [199] and [198] employ FPGA-based hardware-software co-design to enhance the performance of event cameras on drones, highlighting their potential in this field. These approaches leverage the reconfigurability of FPGAs to implement efficient event processing pipelines tailored for drone-specific tasks such as navigation and obstacle avoidance. Additionally, [200] introduced a neuromorphic approach that leverages the Intel Loihi chip to accelerate event camera-driven drones control. By directly feeding event camera data into Spiking Neural Networks (SNNs) running on Loihi, this method enables end-to-end event-driven visual control, significantly improving processing efficiency and responsiveness while capitalizing on Loihi’s inherent low-power neuromorphic architecture.
These major research directions highlight the increasing synergy between specialized hardware and event-based vision. As real-time, edge-based applications such as autonomous driving, robotics, and augmented reality continue to demand low-power, high-speed solutions, these advancements will be crucial in unlocking the full potential of event-based vision in practical deployment scenarios.
Novel event camera sensor designs. Complementing hardware acceleration efforts, novel event camera sensor designs are addressing fundamental limitations while enabling new capabilities, driven by the need to better align with event data characteristics and application demands. These advancements not only address inherent limitations of traditional event cameras but also open new possibilities for their deployment in dynamic and resource-constrained environments.
Significant advancements have been made in event camera sensor technology, as demonstrated in [201] and [202]. These sensors offer higher resolution, reduced pixel pitch, and lower noise-intensity readout modes. Furthermore, [203] introduced a novel generalized event camera that enhances performance in HDR scenes while reducing bandwidth requirements. Its core innovation lies in leveraging Single-Photon Avalanche Diode (SPAD) sensors, which provide high temporal resolution and low readout noise while preserving scene intensity information through adaptive integration. Additionally, lightweight on-chip computation and multi-configuration capabilities further minimize bandwidth usage and power consumption. In [204], a novel event camera sensor design inspired by human microsaccades was introduced, implementing an artificial microsaccade mechanism using a rotating wedge prism to achieve stable texture perception and improved data association, addressing the limitations of motion dependency of traditional event cameras.
Figure 8: Specific hardware and software design for acceleration.
V-B Software acceleration
Motivation. In the deployment of event cameras on mobile platforms, which generate inherently sparse and asynchronous data streams, software acceleration is crucial for efficiently managing and processing high-temporal-resolution, event-driven data in real-time without overburdening resource-constrained devices. Software acceleration plays a pivotal role in optimizing various stages of the event-based processing pipeline, ranging from event denoising and filtering to feature extraction and high-level inference tasks. Specifically, the software acceleration layer transforms raw event streams into structured intermediate representations using event clustering, compression, and sparse matrix manipulation. These optimized representations are then fed into machine learning models or classical computer vision algorithms. Effective software acceleration ensures that data sparsity is fully exploited, computational redundancy is minimized, and real-time performance is maintained, all while adhering to power constraints of mobile applications. The incorporation of sparse data structures, adaptive sampling strategies, and dynamic memory management further underscores the importance of software optimization in unlocking the full potential of event cameras in real-world deployments.
Challenge. Effective software acceleration for event cameras on mobile platforms faces several key challenges that need to be addressed to unlock their full potential. First, the inherent sparsity and asynchronicity of event data pose difficulties for traditional algorithms, which are typically designed for dense inputs. Inefficient processing leads to wasted computational resources and increased latency, necessitating novel algorithms optimized for event-based inputs that can adapt to their unique characteristics. Second, integrating event-based processing within existing deep learning frameworks presents compatibility challenges. Most state-of-the-art neural networks are designed for frame-based data, requiring specialized training pipelines and inference engines to accommodate event representations without sacrificing accuracy or speed. Third, mobile platforms’ limited computational resources further complicate real-time event processing. Efficient memory management and data throughput are critical to balancing processing demands with power constraints. Optimized software must minimize memory usage and reduce data transfer overhead to sustain real-time performance under high workloads. Lastly, existing benchmarking methodologies are largely tailored to conventional imaging systems and fail to capture the unique characteristics of event-driven applications. Developing robust evaluation criteria specific to event-based vision is essential for assessing and improving software acceleration techniques.
Literature review. In the field of event-based vision systems, software acceleration is crucial for enhancing the efficiency and performance of the event processing pipeline. By optimizing specific stages of event handling, researchers can significantly improve computational performance, reduce latency, and optimize resource utilization. Software acceleration can be applied to various stages of the pipeline, including event sampling, preprocessing, feature extraction, event analysis, decision making, and result output. Each stage of the pipeline faces unique challenges and several software acceleration methods have been proposed for these stages:
$\bullet$ Event sampling: Event cameras capture raw event data by recording timestamps and spatial locations. Traditional methods use uniform sampling, but adaptive sampling techniques focus on relevant scene changes. The challenge is balancing temporal resolution and data sparsity. A novel adaptive sampling module using recurrent convolutional SNNs has been proposed in [205], dynamically adjusting the sampling rate based on the spatio-temporal characteristics of events to improve efficiency.
$\bullet$ Event pre-processing: Events need to be denoised, filtered, and formatted to improve subsequent processing. Common challenges include false positives and inefficient filtering. Enhanced filtering algorithms, such as adaptive median filters, can retain important features while reducing computational overhead. A lightweight, hardware-friendly neural network architecture, 2-D CNN, is introduced in [179] for DVS gesture recognition, using a customized median filter to enhance signal-to-noise ratio and reduce hardware complexity.
$\bullet$ Feature extraction: Useful features, including spatial, temporal, and frequency features, are extracted from preprocessed data. Traditional techniques may not effectively capture event data’s unique properties. Software acceleration, particularly through deep learning models tailored for sparse data, can optimize feature extraction while reducing processing time. The FARSE-CNN model, proposed in [206], integrates sparse convolutional and asynchronous LSTM modules for efficient event data processing. [207] introduces SWformer, an attention-free architecture leveraging sparse wavelet transforms to capture high-frequency patterns, resulting in improved energy efficiency and performance.
$\bullet$ Event analysis: This step involves pattern recognition, classification, or regression, commonly for tasks like object detection or tracking. The sparsity of event data can lead to high computational costs. Optimized CNNs and SNNs can enhance the efficiency and effectiveness of analysis. A fast linear solver for camera motion restoration, developed in [208], addresses geometric problems and adapts to sudden motion changes, providing a robust solution for event data.
$\bullet$ Decision making: After event analysis, decisions are made, such as generating control instructions or planning actions. Rapid, accurate decision-making is crucial, and acceleration methods like model compression and fast inference engines can achieve this with lower resource requirements. The TimeLens-XL algorithm, proposed in [209], accurately recovers large motion between frames during video frame interpolation, achieving significant computational cost and model size reduction while ensuring real-time performance.
$\bullet$ Result output: The final stage involves presenting or passing results to downstream systems, focusing on minimizing latency. Software acceleration can streamline data visualization and communication protocols, improving the output phase.
Moreover, existing methods further enhance event processing by optimizing event representation, incorporating geometric and probabilistic models, and leveraging deep learning techniques. Additionally, they exploit the asynchronous and sparse nature of events to accelerate processing while maintaining high accuracy, enabling real-time performance in resource-constrained environments.
Efficient event data representation. Efficient event data representation methods aim to convert the sparse, asynchronous event streams generated by event cameras into a structured format that facilitates efficient processing while preserving critical information and minimizing computational overhead. Recent research has introduced novel event representation techniques, leading to significant advancements in this domain. Notable approaches include event stacking techniques [210], the Temporal Activity Focus (TAF) representation [211], and the HyperHistogram (HH) representation [212]. The TAF method [211] enables adaptive adjustment of time window length and temporal resolution based on spatial locations and event polarity, enhancing the flexibility of event stream representation. Meanwhile, the HH method [211] employs statistical analysis of event polarity and temporal characteristics to construct multiple histograms, which are then integrated into a three-dimensional tensor. This approach effectively preserves polarity information and fine-grained temporal details, addressing the information loss issues prevalent in existing representation techniques.
Geometric models-based methods. Geometric models exploit the spatial relationships between objects and scenes to enhance visual algorithm performance. By incorporating geometric constraints derived from event data, these models mitigate error accumulation, thereby improving estimation accuracy and efficiency. In recent years, significant advancements have been made in event-based algorithms leveraging geometric models, particularly in visual odometry [213], motion estimation [214] [215], and time-to-collision estimation [216], further refining the precision of these tasks.
Probabilistic models-based methods. Probabilistic models provide a mathematical framework that leverages probability distributions to represent data and quantify uncertainty. Given that event camera data is inherently noisy and uncertain, probabilistic models offer an effective means of capturing these characteristics. By employing inference and optimization techniques, these models can enhance the accuracy of event-based vision tasks. In recent years, probabilistic models have been explored in various event-based applications. In object classification, researchers have utilized probabilistic frameworks to maximize the likelihood of a sample belonging to a specific class, effectively leveraging the event stream for learning and improving classification accuracy [217]. For event-based optical flow estimation, probabilistic models have been applied to characterize noise in event data, using belief propagation algorithms for inference and optimization, thereby significantly enhancing estimation accuracy [218].
Deep learning-based methods. Deep learning enables the extraction of complex features from data for prediction and decision-making. Its strong feature extraction and generalization capabilities allow it to learn intricate spatio-temporal patterns from event camera data, leading to more accurate predictions and decisions, thereby enhancing the performance of event-based vision algorithms. In recent years, deep learning-based approaches for event cameras have achieved notable advancements in object detection, motion estimation and tracking, as well as scene understanding and reconstruction. For instance, researchers have explored lightweight networks [211] and sparse Transformer architectures [219] to improve the accuracy and efficiency of event-based object detection. In motion estimation and tracking, deep learning networks have been leveraged for dynamic obstacle avoidance [220], while the integration of motion cues and object contour information has enabled robust object tracking [221]. For scene understanding and reconstruction, dynamic graph convolutional networks (GCNs) and cross-representation distillation frameworks have been employed to enhance event-based visual learning [222]. Additionally, researchers have developed point cloud-based architectures to improve pose relocalization accuracy [141] and have utilized a unified implicit neural representation framework to enhance rolling shutter image restoration [223].
SNNs are computational models that emulate the functioning of biological neurons, using spike-based signals to transmit information. They offer key advantages such as low power consumption and high parallelism, making them particularly well-suited for processing the asynchronous and sparse nature of event camera data. By leveraging these properties, SNNs enable more efficient and adaptive event-based vision algorithms. In recent years, significant progress has been made in utilizing SNNs for efficient event-driven computation across various domains, including object classification, optical flow estimation, drone control, video frame interpolation, sparse recurrent SNN simulation and training, and novel SNN architecture design. For instance, researchers have explored SNN-based object classification [217, 224], integrated SNNs with artificial neural networks (ANNs) for optical flow estimation [225], and employed SNNs for drone landing control [226]. Additionally, SNNs have been leveraged for high-speed content reconstruction [227], the development of event-driven algorithms for simulating and training sparse recurrent SNNs [228], and the design of novel SNN architectures [229].
Asynchronous and sparse computation-based acceleration. Asynchronous and sparse computation involves performing calculations only when data changes, processing only the modified components. This approach capitalizes on the inherent asynchronous and sparse nature of event camera data, significantly reducing computational costs and improving the efficiency of event-based vision algorithms. For example, algorithms can be designed to perform computations only when pixel brightness changes, or to extract and analyze features solely in areas where such changes occur, thereby optimizing computational efficiency. Recent advancements have been made in accelerating event camera vision processing through asynchronous and sparse computation techniques. Notable efforts include the development of network architectures for efficiently processing asynchronous event streams [230], as well as the creation of graph-based frameworks for asynchronous event processing [231]. Additionally, researchers have explored the use of local shift operations and specialized asynchronous event processing mechanisms to enhance the efficiency of event-based processing [232].
These advancements in software acceleration significantly improve the efficiency and applicability of event-based vision systems, particularly for real-time mobile applications.
VI Application: Mobile platform-based task
This section explores the various tasks associated with mobile platforms, emphasizing the use of event cameras and other sensing technologies to improve performance in dynamic environments. These tasks are categorized into intrinsic sensing, external sensing, and event-based SLAM, each offering distinct applications and advantages, as shown in Fig.9.
Figure 9: Event camera-based sensing tasks on mobile platforms
VI-A Intrinsic sensing
Intrinsic sensing involves the use of internal sensors on mobile platforms to collect data about the immediate environment. This encompasses visual odometry, and optical flow, with event cameras providing high temporal resolution and low-latency data that enhance accuracy and efficiency.
Vision odometry. Vision odometry uses event cameras to estimate displacement during high-speed motion, offering greater precision and lower computational demands compared to traditional methods[233],[213],[234]. Current approaches for pose or motion parameter estimation primarily rely on information directly extracted from the event stream. A notable study in this area, [235], focuses on angular velocity estimation. Its key innovation lies in an algorithm based on contrast maximization, which effectively leverages edge information from the event stream. Another significant work enhances attitude estimation accuracy and efficiency in high-rotation scenarios through improved aggregation functions [236]. Additionally, [237] proposes DEVO, a visual odometry system that integrates depth and event camera data and is designed to operate in challenging conditions. The first event-based stereo visual-inertial odometry system is introduced in [238], integrating stereo event cameras, standard cameras, and IMUs, and employing spatiotemporal correlation and motion compensation techniques to further improve robustness and accuracy in dynamic environments.
Optical flow. Event cameras are highly effective in optical flow calculations, enabling real-time processing of high dynamic range scenes essential for accurate motion tracking. Optical flow estimation methods based on event cameras can be broadly classified into two types: deep learning-based techniques and traditional computer vision-based algorithms.
Deep learning-based methods. These methods exploit the feature extraction capabilities of deep neural networks to estimate optical flow directly from the event stream. For example, [239] proposed a self-supervised learning approach that does not require ground truth optical flow annotations, thereby reducing data labeling costs. Additionally, optical flow estimation via a hierarchical spiking neural network was demonstrated in [240], enabling global motion perception. These methods often require preprocessing or transforming event data into a suitable format for deep learning models.
Traditional computer vision-based methods. These methods directly utilize the spatio-temporal information from the event stream for optical flow calculations. An event-based method for normal flow calculation was introduced in [241], which, when combined with IMU measurements, facilitates fusion within a continuous time framework for linear velocity estimation. This method enhances the direct processing of event data, making it especially suited for high dynamic motion scenarios. Furthermore, the use of contrast maximization methods for optical flow estimation was explored in [242], along with an in-depth study of reward functions. The value of optical flow estimation in robotic control was demonstrated in [243], which utilized event camera flow information for drone heading estimation.
VI-B External sensing
External sensing refers to the use of sensors to gather data about the broader environment surrounding the mobile platform. This encompasses a range of tasks, including mapping, object detection and tracking, obstacle avoidance, optical flow, classification, 3D reconstruction, and segmentation, all of which benefit from the high temporal resolution and low latency provided by event cameras.
Mapping. Event cameras enhance the accuracy and efficiency of 3D map construction by capturing rapid environmental changes. Feature-based methods are commonly employed for map building, where the primary goal is the extraction of feature points or line features from the scene. For example, straight-line features in railway scenes are detected using an event camera in [244], which combined odometry information for triangulation and map construction of railway infrastructure. However, such methods rely on the stability and repeatability of features, which can be problematic in environments with sparse feature points or significant occlusions.
To address these challenges, novel methods have been proposed. BeNeRF allows the recovery of a neural radiance field (NeRF) from a single blurred image and its corresponding event stream [245]. This approach does not require extensive camera pose data, and the learned scene representation can be applied for subsequent map construction. Additionally, the AsynHDR system [246] captures high dynamic range (HDR) scenes and reconstructs HDR images by modulating incident light through an LCD panel. The generated HDR images provide richer information for 3D reconstruction.
Object detection & tracking. The high temporal resolution and low latency of event cameras make them particularly well-suited for precise object detection and tracking in dynamic environments[247],[248],[249],[250],[219],[251]. These methods can be broadly classified into two categories: event stream-based detection and tracking, and fusion-based detection and tracking that integrates additional information sources (such as frame images or IMU data) to enhance performance.
Event stream-based detection and tracking methods. These methods rely primarily on the spatio-temporal characteristics of the event stream for object detection and tracking. For instance, a novel stereo event-based tracking algorithm was proposed in [252], which addresses object occlusion by combining 3D reconstruction with cluster tracking. Furthermore, EventCap [253] and EventEgo3D [254] utilize monocular event streams for capturing 3D human motion, highlighting the potential of event cameras in human pose estimation and tracking. In [255], corner tracking is employed to filter noise and inconsistent corners, and APM is used to identify events generated by moving objects (intruders), ultimately tracking intruders through clustering high-priority events. EVPropNet [256] tracks drone propellers using event data, while the EDOPT [257] relies solely on event cameras for six degrees of freedom object pose tracking.
Fusion-based detection and tracking methods. These methods integrate additional data sources to enhance the accuracy and robustness of detection and tracking. In [258] and [259], frame images and event data are fused to improve target detection. The combination of event streams and video frames for person re-identification is explored in [260]. In [3], mmWave radar and event stream are combined to localize the drone with high frequency. Meanwhile, in [261], depth camera and event stream are integrated to track the obstacle.
Optical Flow. Event cameras excel in external optical flow calculations, enabling accurate motion tracking in high-speed scenarios [262],[263],[264],[265],[218]. Unlike internal perception, which focuses on the camera’s motion, optical flow information here is used to understand and predict the movement of external objects. Typical applications include high-speed object capture and robotic control.
RPEFlow[266] enhance accuracy by combining event data with RGB and point clouds, leveraging cross-modal attention for dynamic scenes. Contrast Maximization (CM) based methods[267][268] offer a principled alternative, achieving state-of-the-art performance through innovations like multi-reference focus loss and time-aware flow modeling. Distinct from these, TEGBP[218] provides an efficient, incremental estimation of full optical flow from sparse event data. Collectively, these advancements demonstrate the diverse and potent strategies emerging to harness event cameras for robust external motion analysis.
Classification. Event cameras’ ability to capture rapid changes in lighting and motion sigOnenificantly improves the performance of image classification tasks[269],[270],[271]. Classification methods based on event cameras leverage the subtle motion and lighting changes captured by the cameras, often combining these with various models for recognition. One study [272] converts the event stream into a graph structure and applies graph convolutional networks for classification, effectively modeling the spatiotemporal relationships in the data. In [273], a deep learning model for human gait recognition was used to distinguish individuals by learning motion patterns from the event stream. Additionally, the classification of micro-particles using an event camera and SNNs was demonstrated in [274], highlighting the potential of event cameras for microscopic object recognition. In [275], facial micro-expressions are captured using an event camera for emotion classification, demonstrating the value of event cameras in recognizing subtle expressions.
3D Reconstruction. Event cameras facilitate high-precision 3D reconstruction in dynamic environments, leveraging their high temporal resolution and sparse event stream. Real-time 3D reconstruction using a monocular event camera was demonstrated in [276], highlighting the potential of event cameras for such tasks. In [277], semi-dense 3D reconstruction was performed using a stereo event camera, achieving high accuracy while maintaining speed. The integration of structured light technology with event cameras enabled high-speed 3D scanning in [278], benefitting from the low data redundancy of event cameras. Furthermore, a method for 3D shape reconstruction based on polarization information [279] maintains high precision even at low event rates.
Several methods focus on reconstructing specific objects or scenes. For example, EventHands [280] employed event streams for real-time 3D gesture estimation, while EventHPE [281] estimated 3D human pose and shape from event data, demonstrating event cameras’ capability in capturing human posture. Additionally, [282] focused on the reconstruction of non-rigid object deformations, highlighting the ability of event cameras to handle complex motion scenarios.
Segmentation. Event cameras enhance image segmentation performance by providing accurate and efficient segmentation in rapidly changing environments[283],[284],[285]. Segmentation methods based on event cameras can be categorized into motion-based segmentation and deep learning-based semantic segmentation.
Motion-based segmentation methods utilize motion information captured by event cameras to segment images. For instance, a motion segmentation method based on motion compensation was proposed in [286], which uses an iterative optimization algorithm to segment the scene into independently moving objects. This method effectively exploits the high sensitivity of event cameras to motion changes, enabling dynamic target segmentation.
Deep learning-based semantic segmentation methods take advantage of the powerful feature extraction and semantic understanding capabilities of deep neural networks. The CMDA method [287] used the high dynamic range characteristics of event cameras to compensate for the limited dynamic range of traditional cameras in nighttime scenes, achieving robust semantic segmentation in such conditions.
VI-C Simultaneous Localization and Mapping (SLAM)
Event-based Simultaneous Localization and Mapping leverages the inherent advantages of event cameras—high temporal resolution and low latency—to achieve robust and accurate localization and mapping, particularly within challenging, dynamic environments[288],[172],[289],[290]. The speed and accuracy demonstrated by event-based SLAM algorithms highlight the transformative potential of these sensors in revolutionizing SLAM applications.
A prominent strategy to enhance robustness is multimodal sensor fusion, where the limitations of individual sensors are mitigated by integrating complementary data streams. Ultimate SLAM [291] exemplifies this trend, achieving robust visual SLAM in high dynamic range and high-speed scenarios through a tightly coupled fusion of event, standard frame camera, and IMU data. Similarly, Implicit Event-RGBD Neural SLAM [292] demonstrates improved robustness in non-ideal conditions like motion blur and illumination changes by fusing event and RGB-D data within an implicit neural representation framework. Beyond sensor fusion, innovations in feature representation are also crucial. Event-based Line SLAM in Real-time, for example, utilizes line features to address the feature loss challenges associated with point-based methods in fast motion or low-light conditions, improving efficiency through a Parallel Tracking and Mapping (PTAM) framework.[293] Furthermore, recent advancements explore optimization-based approaches, with CMax-SLAM (2024) representing a significant contribution.[294] This method introduces a Contrast Maximization (CMax) based rotational motion bundle adjustment, directly leveraging event data for global optimization and enhancing the precision of rotational motion estimation. These diverse advancements underscore the rapid progress and multifaceted approaches within the evolving field of event-based SLAM.
VII Future direction and discussion
Although there has been extensive research on event cameras, they are still in the early stages of development for mobile sensing. Significant opportunities remain for advancement in both research and industry, with several trade-offs to consider, such as latency, power consumption, and accuracy. Below, we outline key future directions for event-based mobile sensing.
(1) Improving event cameras using optics devices. Event cameras generate responses only when changes in illumination intensity occur within their field of view. As a result, they produce no output in completely static scenes, limiting their applicability in scenarios where continuous perception is required. Future research could explore integrating optical elements to enhance event cameras’ responsiveness under static or low-dynamic conditions. For instance, dynamic optical components, such as tunable lenses or light modulators, could introduce controlled variations in illumination, enabling event cameras to perceive otherwise static environments. Additionally, hybrid optical-electronic solutions, such as coupling event cameras with active illumination sources or computational imaging techniques, may further extend their capabilities in challenging visual conditions. These advancements could significantly improve event cameras’ adaptability for tasks requiring persistent environmental awareness, such as long-term surveillance, augmented reality, and autonomous navigation.
(2) Designing neuromorphic hardware to facilitate event processing. Current hardware exhibits significant inefficiencies in handling event-based processing tasks. When event-based systems run on conventional CPUs, frequent context switching disrupts event processing, leading to substantial drops in event throughput. While GPUs excel in managing traditional vision workloads, they are not inherently optimized for handling asynchronous, high-frequency event streams, resulting in suboptimal performance. FPGAs offer improvements in event processing by enabling parallelism and low-latency operations. However, they lack comprehensive pipeline optimization, limiting their efficiency in end-to-end event-based vision tasks. To address these challenges, future research should focus on developing specialized neuromorphic hardware tailored for event-driven computation. Dedicated neuromorphic chips, such as those leveraging SNNs or asynchronous processing architectures, could significantly enhance efficiency by aligning with the native characteristics of event data. Additionally, designing hybrid architectures that integrate neuromorphic computing with conventional accelerators (e.g., FPGA-GPU co-processing frameworks) could further improve scalability and adaptability across various event-based applications, including real-time perception for autonomous systems, high-speed robotics, and ultra-low-power edge computing.
(3) Leveraging the complementary strengths of event camera and other sensors. Event cameras exhibit complementary characteristics with various other sensing modalities, such as frame-based cameras, mmWave radar, and LiDAR. Integrating these diverse sensors can enhance the robustness, accuracy, and applicability of event-based vision systems across a wide range of scenarios. Frame cameras provide rich texture and color information, compensating for event cameras’ inability to capture absolute intensity values in static scenes. Meanwhile, event cameras offer high temporal resolution and low-latency response, overcoming motion blur and improving dynamic scene perception. The fusion of these modalities can enable more accurate visual odometry, object detection, and scene reconstruction. Beyond vision-based sensing, mmWave radar and LiDAR introduce depth and environmental perception capabilities that further complement event cameras. mmWave radar is robust to adverse lighting and weather conditions, making it a valuable addition for applications such as autonomous navigation and human activity recognition. LiDAR, on the other hand, provides precise 3D structural information, which can be combined with the fine-grained temporal data from event cameras to enhance 3D reconstruction, depth estimation, and SLAM. Future research should explore more effective sensor fusion strategies, including deep learning-based multi-modal integration, cross-domain feature alignment, and adaptive fusion frameworks. Additionally, optimizing hardware architectures for real-time, low-power sensor fusion remains a crucial challenge for deploying event-based multi-modal sensing systems in mobile robotics, autonomous vehicles, and wearable devices.
(4) Bio-inspired algorithm design. Event cameras, modeled after the human retina, inherently exhibit neuromorphic characteristics, making them well-suited for bio-inspired algorithm design. By drawing inspiration from biological vision systems, researchers can develop more efficient and adaptive computational models for event-based perception. One promising direction is the integration of SNNs, which process information through discrete spikes rather than continuous signals, closely mimicking the neural activity of the brain. SNNs offer advantages such as sparse computation, low power consumption, and high temporal precision, making them particularly suitable for event-driven vision tasks like object recognition, optical flow estimation, and autonomous navigation. Beyond SNNs, bio-inspired motion perception models, such as those based on the primate visual cortex, can be leveraged to improve event-based tracking and motion segmentation. Techniques such as hierarchical processing and attention mechanisms can enhance the system’s ability to extract meaningful spatio-temporal patterns from sparse and asynchronous event streams. Additionally, neuromorphic control systems, inspired by biological sensorimotor loops, can enhance real-time decision-making in robotics and autonomous systems. By embedding event-based feedback mechanisms and bio-plausible learning rules, such systems could achieve faster reactions and more robust performance in dynamic environments. Future research should explore hybrid architectures that combine bio-inspired principles with modern deep learning techniques, enabling more biologically plausible yet computationally efficient event-based vision algorithms. Moreover, optimizing neuromorphic hardware to support such models remains an essential step toward the deployment of bio-inspired event processing systems.
(5) Hardware-software co-optimization techniques. The efficient processing of event data requires a synergistic integration of hardware and software. Current hardware architectures often struggle with the unique characteristics of event-based vision, leading to suboptimal performance when traditional computing paradigms are applied. To address these challenges, hardware-software co-optimization techniques are essential for unlocking the full potential of event-based systems. One promising direction is the co-design of dedicated neuromorphic hardware tailored for event processing. This includes the development of custom accelerators, such as SNN chips, FPGA-based event processors, and optimized ASIC architectures. These specialized hardware solutions can significantly enhance efficiency by directly supporting event-driven computations, reducing the overhead associated with conventional CPUs and GPUs. At the software level, event-driven programming models and efficient data structures are needed to minimize latency and computational redundancy. Traditional deep learning frameworks, designed for dense frame-based input, are often inefficient for event-based data. Thus, lightweight and adaptive architectures, such as event-driven neural networks, graph-based processing frameworks, and sparse computation techniques, must be further explored to better align with hardware capabilities. Furthermore, cross-layer optimizations—where algorithms, runtime systems, and hardware work in concert—can lead to significant performance improvements. For example, dynamically allocating computational resources based on event activity, optimizing memory access patterns to reduce bottlenecks, and leveraging hardware-aware model compression techniques can enhance both speed and energy efficiency. Future research should focus on holistic co-design methodologies that bridge the gap between event-based algorithms and emerging neuromorphic hardware. By refining the interplay between hardware and software, event-driven vision systems can achieve real-time performance with minimal power consumption, making them more viable for applications in robotics, autonomous vehicles, and edge computing.
VIII Conclusion
Event-based vision has emerged as a transformative paradigm for mobile sensing, offering high temporal resolution, low latency, and energy efficiency. This survey has provided a comprehensive overview of event cameras, covering their principles, data representation techniques, algorithmic developments, hardware and software acceleration strategies, and diverse applications in mobile platforms. Despite their advantages, event cameras face several challenges, including effective event processing, sensor fusion, and real-time deployment on resource-constrained platforms. Future research directions should focus on improving event camera hardware with advanced optics, designing neuromorphic processors tailored for asynchronous processing, and leveraging bio-inspired algorithms to enhance perception capabilities. Additionally, integrating event cameras with complementary sensors, such as LiDAR and radar, will further expand their applicability in dynamic environments. We hope this survey serves as a valuable reference, inspiring further research and facilitating the practical deployment of event-based vision systems in mobile sensing community.